OpenStudy (anonymous):

Which of the following are the x-intercepts on the graph of the function shown below? f(x)=2(x+1)(x-2) More than one answer A. 2 B. -1 C. -4 D. -2 Which of the following are the x-intercepts on the graph of the function shown below? f(x)=2(x+1)(x-2) More than one answer A. 2 B. -1 C. -4 D. -2 @Mathematics

jimthompson5910 (jim_thompson5910):

f(x)=2(x+1)(x-2) 0=2(x+1)(x-2) 2(x+1)(x-2)=0 x+1=0 or x-2=0 x=-1 or x=2

jimthompson5910 (jim_thompson5910):

So the answers are choices A and B

OpenStudy (anonymous):

A and B